OpenStudy (anonymous):

Which of the following is equivalent to the proportion 2/3 =3y? A. 3y = (2)(3) B. 2y = (2)(3) C. 2y = (3)(3) D. 3y = (3)(3)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not really

OpenStudy (skullpatrol):

None are right.

OpenStudy (anonymous):

how is that possible i don't make the question lol

OpenStudy (skullpatrol):

Check this 2/3 =3y is it suppose to be 2/3 =3/y?

OpenStudy (skullpatrol):

If so then C. 2y = (3)(3)

OpenStudy (anonymous):

the Y supposed to = 9?

OpenStudy (skullpatrol):

Nope, y = 9/2